OpenStudy (daenio):

What is the concentration of H+ in a solution containing 73.0g of HCL in 500 mL of water?

OpenStudy (anonymous):

H+ = 73/36.5 * 1 = 2gm of H+ no of moles = 2/1 = 2 Concentration (Molarity (M) = \[\frac{ 2*1000 }{ 500 } = 4M\]

OpenStudy (daenio):

I actually figured this one out. Thanks for confirming the solution! :D

OpenStudy (anonymous):

Welcome